# The Growing Gains and Pains of Iterative Web Corpora Crawling: Insights from South Slavic CLASSLA-web 2.0 Corpora

### v1

@@ 1.  Introduction
Our study makes the following contributions. First, we release a substantially larger web corpus collection containing 17.0 billion words in 38.1 million texts across Bosnian, Bulgarian, Croatian, Macedonian, Montenegrin, Serbian, and Slovenian (presented in Section 4). The CLASSLA-web 2.0 corpora can be downloaded in JSONL and VERT formats from the CLARIN.SI repository (Kuzman Pungeršek et al., 2026) [^1] The CLASSLA-web 2.0 corpora are freely accessible on the CLARIN.SI repository at http://hdl.handle.net/11356/2079. and queried through the CLARIN.SI concordancers [^2] https://www.clarin.si/ske/, the links to each of the corpora are provided at the website https://clarinsi.github.io/classla-web/.. Each text is linguistically annotated and labelled with genre and topic categories. Leveraging these annotation layers, we provide a detailed view of genre and news topic distributions across South Slavic webs (Section 4.1).
@@ 2.  Related Work
The practice of building large web corpora for European languages was established with the WaCky initiative (Baroni et al., 2009). The first web corpora of South Slavic languages were compiled for Slovenian (slWaC) and Croatian (hrWaC) (Ljubešić and Erjavec, 2011; Erjavec et al., 2015), followed by Bosnian (bsWaC) and Serbian (srWaC) (Ljubešić and Klubička, 2014). After a seven-year gap, the MaCoCu (Bañón et al., 2022) project resumed crawling activities for South Slavic languages and other less-resourced European languages. The monolingual web corpora for South Slavic languages that were produced in the MaCoCu project have been later additionally post-processed and enriched with linguistic and genre automatic annotation, and published as CLASSLA-web 1.0 (Ljubešić and Kuzman, 2024) corpus collection. The CLASSLA-web corpora have received considerable attention in the South Slavic space, both from the NLP community (Ljubešić et al., 2024b; Vreš et al., 2024; Kuzman and Ljubešić, 2025; Kuzman et al., 2024) and from linguists and language teachers (Erjavec et al., 2024; Ljubešić et al., 2024a).
@@
In parallel, numerous other web-based datasets have emerged from the Common Crawl and the Internet Archive data collections, e.g., the cc100 dataset (Conneau et al., 2020), the mC4 dataset (Xue et al., 2021), the OSCAR dataset (Suárez et al., 2019), and most recent, the HPLT (de Gibert et al., 2024) and FineWeb2 datasets (Penedo et al., 2025). While these multilingual collections are highly useful for multilingual language modelling tasks, they suffer, inter alia, from lower quality of content, namely, cc100 and mc4 datasets (van Noord et al., 2024); an unexplainable small amount of content in certain South Slavic languages, namely in OSCAR [^3] https://oscar-project.github.io/documentation/versions/oscar-2301/, mC4 dataset (Xue et al., 2021), the HPLT 3.0 datasets (de Gibert et al., 2024), and the FineWeb2 dataset (Penedo et al., 2025). More specifically, the most recent HPLT 3.0 (de Gibert et al., 2024) and FineWeb2 (Penedo et al., 2025) datasets are missing Montenegrin language. Moreover, they report unusual sizes of Serbian and Bosnian corpora where the Bosnian web corpus is significantly larger than the Serbian one, which does not correlate with the number of language speakers or web corpus sizes reported for CLASSLA-web corpora (Ljubešić and Kuzman, 2024). It seems that these corpora that use “one pipeline to scale them all” suffer from language identification issues related to the Serbian, Bosnian and Montenegrin languages. Furthermore, since the introduction of large language models that make text generation in various languages easily available, there is a growing problem on the web that more and more content is automatically generated, as we will show in Section 5.3. It is thus crucial that manual inspection of web corpora content by native speakers is included in the corpora preparation to filter out as much bad content as possible, and to assure high quality of data. Our work addresses these gaps via targeted TLD crawling, specialized language discrimination between the mutually intelligible Croatian, Serbian, Bosnian and Montenegrin languages, and manual domain validation.
@@ 3.  CLASSLA-web Construction Pipeline
The goal of the crawling pipeline is to produce large, clean, and richly annotated web corpora. As shown in Figure 1, the process consists of 1) crawling based on top-level national domains, and post-processing, 2) automatic text annotation with genre and topic information, and 3) linguistic annotation. In this section, we briefly present each step. For more details, refer also to the description of the CLASSLA-web crawling pipeline in Ljubešić and Kuzman (2024) and to the website presenting the CLASSLA-web collections (https://clarinsi.github.io/classla-web/).
@@
Following the methodology set up in the MaCoCu project (Bañón et al., 2022), we crawl national top-level domains (e.g., .si for Slovenian), as well as connected generic domains (.com and others). Crawling is performed with the MaCoCu crawler [^4] https://github.com/macocu/MaCoCu-crawler, which is based on the SpiderLing crawler (Suchomel et al., 2012) and designed for efficient large-scale web text collection.
@@
Two language identification tools are used at document and paragraph levels, namely, a trigram classifier (Lui and Baldwin, 2014) and the Compact Language Detector 2 (CLD2) [^5] https://github.com/CLD2Owners/cld2 (Sites, 2013). Language identification between Bosnian, Croatian, Montenegrin, and Serbian, comprising the HBS macro-language, has been shown to be particularly challenging as these South Slavic languages exhibit a significant level of mutual intelligibility (Rupnik et al., 2023). We directly assign texts that come from national top-level domains to the corresponding web corpus, e.g., web domains from .hr are included in the Croatian web corpus. For generic top-level domains, we use an additional language identification tool specialized for distinguishing between Bosnian, Croatian, Montenegrin and Serbian, namely a Naive Bayes classifier trained on language-specific wordlists extracted from national TLDs (Rupnik et al., 2023). During the development of CLASSLA-web 2.0 corpora, we also explored language identification through zero-shot prompting with large language models, but the specialized classifier proved to be more reliable for HBS disambiguation.
@@
After crawling, we remove boilerplate with the jusText [^6] http://corpus.tools/wiki/Justext tool (Pomikálek, 2011) and near-duplicates with the onion [^7] http://corpus.tools/wiki/Onion tool (Pomikálek, 2011), and discard exact duplicates. We discard texts that are not written in the target language, repair encoding, clean unwanted HTML elements, and filter overly short documents (comprising of only short paragraphs with less than 70 characters, or consisting of less than 75 words).
@@
In CLASSLA-web 2.0 corpora, a new annotation layer is introduced – in addition to genres, texts are annotated with topics to provide an additional insight into the corpora content. We use a multilingual news topic classifier (Kuzman and Ljubešić, 2025) covering 17 top level labels from the IPTC Media Topic NewsCodes schema (e.g., Politics, Science and Technology, Sport – for a detailed description of labels, refer to Kuzman and Ljubešić (2025)). On a manually-annotated test set in Croatian, Slovenian, Catalan, and Greek, the model achieves macro-F1 of 0.75 and micro-F1 of 0.73. For instances assigned with the class probabilities below 0.6, we use the Mix label. As with the genre annotation, the confidence threshold was identified based on a manual analysis. We should note that the classifier has been for now evaluated only on news texts, as it was primarily intended for news topic classification. Thus, in Section 4.1, we analyse topics only within the News genre.
@@
All texts are linguistically annotated with the CLASSLA-Stanza (Ljubešić et al., 2024c) [^9] https://pypi.org/project/classla/ pipeline, providing tokenization, lemmatization and morphosyntactic annotation with state-of-the-art accuracy for Slovenian [^10] Based on the SloBENCH benchmark at https://slobench.cjvt.si/leaderboard/view/11., and a high performance on Croatian, Serbian, Bulgarian, and Macedonian (Ljubešić et al., 2024c). We use the “web” module of the pipeline, which has been shown to perform well on heterogeneous web texts. The Croatian module is applied to Bosnian and Montenegrin. A more detailed argumentation behind the module choices is provided in Ljubešić et al. (2024c).
@@
While the overall CLASSLA-web pipeline is more or less the same between versions 1.0 and 2.0 of CLASSLA-web corpora, three improvements were implemented during the development of the CLASSLA-web 2.0 version. First, near-duplicate removal was improved by masking numbers, punctuation, and links during paragraph-level deduplication to improve the efficiency of identifying texts that differ only in numbers. Second, we expanded manual inspection of top domains from only Slovenian and Croatian in CLASSLA-web 1.0 to all languages in CLASSLA-web 2.0. Third, we added topic annotation to complement genre labels.
@@
The CLASSLA-web 2.0 collection comprises all seven South Slavic national webs and languages – Bosnian (bs), Bulgarian (bg), Croatian (hr), Macedonian (mk), Montenegrin (cnr), Serbian (sr) and Slovenian (sl). In total, it consists of 17 billion words and 38 million texts, as reported in Table 1. The biggest corpus, namely, the Bulgarian (CLASSLA-web.bg 2.0) corpus, comprises 6 billion words and 15 million texts. Other large corpora are Serbian (CLASSLA-web.sr 2.0), Croatian (CLASSLA-web.hr 2.0) and Slovenian (CLASSLA-web.sl 2.0) with 2.3 to 3.7 billion words and 5 to 7 million texts. Bosnian (CLASSLA-web.bs 2.0) and Macedonian (CLASSLA-web.mk 2.0) corpora are smaller, comprising 0.7 to 1 billion words and 2.1 to 2.5 million texts. The Montenegrin (CLASSLA-web.cnr 2.0) corpus is the smallest with 0.3 billion words and 0.8 million texts. However, as shown in Section 5.1, the new Montenegrin corpus is twice the size of the Montenegrin corpus included in the CLASSLA-web 1.0 collection and, to the best of our knowledge, represents the largest general Montenegrin corpus available.
@@ 4.1.  Genre and Topic Distribution
Genre and topic labels provide valuable insights into similarities as well as differences between web corpora. As shown in Figure 2, News content dominates in all national webs. This genre is especially dominant in Bosnian, Macedonian, and Montenegrin corpora, where it exceeds 70%, whereas it represents 30–40% of texts in Croatian and Slovenian corpora. As in CLASSLA-web 1.0 corpora (Ljubešić and Kuzman, 2024), Promotion is more prevalent in Slovenian, Croatian and Bulgarian web corpora (12–20%) than in others. Another genre where there is a visible difference between corpora is Forum. Texts from forums seem to be much more frequent in Bulgarian, Serbian, and Slovenian webs (10–13%) than on Bosnian, Macedonian, and Montenegrin webs (below 5%).
@@
As News is the predominant genre in all web corpora and given that the topic classifier has been primarily validated on news texts, we analyse topics within the News genre. In Figure 3, we show the distribution of topics that are discussed in more than 10% of news texts in at least one corpus. Despite the fact that the news topic schema comprises 17 topic labels, only five categories appear with high enough frequency in the seven corpora, namely, “Sport”, “Politics”, “Economy, Business and Finance”, “Arts, Culture, Entertainment and Media”, and “Human Interest”, in this order of average frequency across corpora. These five categories account for approximately 60% of news texts in each corpus. Sport is the most common topic of news texts in the majority of web corpora, and it even represents a fifth of all news in the Macedonian web corpus. In Bulgarian and Montenegrin web corpora, the most frequent topic of news texts is Politics, accounting for 20% of the news texts within the Montenegrin corpus. The distribution of the other three most frequent topic categories is relatively consistent across the South Slavic corpora, except in the case of Arts, Culture, Entertainment and Media which is notably more prevalent in the Croatian corpus compared to the others, representing almost 15% of Croatian news texts.
@@
To understand how national webs evolve over short time spans, we compare CLASSLA-web 2.0 corpora with CLASSLA-web 1.0 (Ljubešić and Kuzman, 2024). Both web corpus collections were compiled using an identical construction pipeline (see Section 3), with the 1.0 version having been collected from the web two to three years prior to the 2.0 version, specifically in 2021 and 2022. In the following sections, we examine the size difference between the two corpora collections (Section 5.1), content overlap (Section 5.2), and the presence of automatically generated texts and other low quality content (Section 5.3).
@@ 5.1.  Size Gains
As shown in Figure 4, the second crawling iteration produced larger web corpora, showing the increase of content on the web. Compared to the size of CLASSLA-web 1.0 corpora, the web corpora in the 2.0 corpus collection comprise 57% more words and 46% more texts, that is, roughly half as many words and texts as in the first version. Interestingly, the web growth is uneven across South Slavic webs – the Slovenian and Croatian web exhibits modest increase of size, with a 30% percent change in size in terms of number of words, while the size of Bulgarian and Montenegrin web corpora nearly doubled in the second crawling iteration. These differences in sizes reflect the high growth of content on the web rather any differences in the crawling methodology, as both CLASSLA-web 1.0 and CLASSLA-web 2.0 corpora have been produced with practically the same pipeline.

### v2

# arXiv:2601.11170v2[cs.CL] 27 Feb 2026

@@ 1.  Introduction
Our study makes the following contributions. First, we release a substantially larger web corpus collection containing 17.0 billion words in 38.1 million texts across Bosnian, Bulgarian, Croatian, Macedonian, Montenegrin, Serbian, and Slovenian (presented in Section 4). The CLASSLA-web 2.0 corpora can be downloaded in JSONL and VERT formats from the CLARIN.SI repository (Kuzman Pungeršek et al., 2026) [^1] The CLASSLA-web 2.0 corpora are freely accessible in the CLARIN.SI repository at http://hdl.handle.net/11356/2079. and queried through the CLARIN.SI concordancers [^2] https://www.clarin.si/ske/, the links to each of the corpora are provided on the website https://clarinsi.github.io/classla-web/.. Each text is linguistically annotated and labelled with genre and topic categories. Leveraging these annotation layers, we provide a detailed view of the distribution of genres and news topics across South Slavic webs (Section 4.1).
@@ 2.  Related Work
The practice of building large web corpora for European languages was established with the WaCky initiative (Baroni et al., 2009). The first web corpora of South Slavic languages were compiled for Slovenian (slWaC) and Croatian (hrWaC, Ljubešić and Erjavec, 2011; Erjavec et al., 2015), followed by Bosnian (bsWaC) and Serbian (srWaC, Ljubešić and Klubička, 2014). After a seven-year gap, the MaCoCu project (Bañón et al., 2022) resumed crawling activities for South Slavic languages and other less-resourced European languages. The monolingual web corpora for South Slavic languages that were produced in the MaCoCu project have later been additionally post-processed and enriched with linguistic and genre annotation, and published as CLASSLA-web 1.0 corpus collection (Ljubešić and Kuzman, 2024). The collection spans Bosnian Ljubešić et al. (2024a), Bulgarian Ljubešić et al. (2024b), Montenegrin Ljubešić et al. (2024e), Croatian Ljubešić et al. (2024c), Macedonian Ljubešić et al. (2024d), Slovenian Ljubešić et al. (2024g), and Serbian Ljubešić et al. (2024f) web corpora. The CLASSLA-web corpora have received considerable attention in the South Slavic space, both from the NLP community (Ljubešić et al., 2024b; Vreš et al., 2024; Kuzman and Ljubešić, 2025; Kuzman et al., 2024) and from linguists and language teachers (Erjavec et al., 2024; Ljubešić et al., 2024a).
@@
In parallel, numerous other web-based datasets have emerged from the Common Crawl and Internet Archive data collections, e.g., the cc100 dataset (Conneau et al., 2020), the mC4 dataset (Xue et al., 2021), the OSCAR dataset (Suárez et al., 2019), and most recent, the HPLT (de Gibert et al., 2024) and FineWeb2 datasets (Penedo et al., 2025). While these multilingual collections are highly useful for multilingual language modelling tasks, they suffer, inter alia, from lower quality of content, namely, cc100 and mc4 datasets (van Noord et al., 2024); or an unexplainable small amount of content in certain South Slavic languages, namely in OSCAR [^3] https://oscar-project.github.io/documentation/versions/oscar-2301/, mC4 dataset (Xue et al., 2021), the HPLT 3.0 datasets (de Gibert et al., 2024), and the FineWeb2 dataset (Penedo et al., 2025). More specifically, the most recent HPLT 3.0 (de Gibert et al., 2024) and FineWeb2 (Penedo et al., 2025) datasets are missing Montenegrin language. Moreover, they report unusual sizes of Serbian and Bosnian corpora where the Bosnian web corpus is significantly larger than the Serbian one, which does not correlate with the number of language speakers or web corpus sizes reported for the CLASSLA-web corpora (Ljubešić and Kuzman, 2024). It seems that these corpora that use “one pipeline to scale them all” suffer from language identification issues related to the Serbian, Bosnian and Montenegrin languages. Furthermore, since the introduction of large language models that make text generation in various languages easily available, there is a growing problem on the web that more and more content is automatically generated, as we will show in Section 5.3. It is thus crucial that manual inspection of web corpora content by native speakers is included in the corpora preparation to filter out as much bad content as possible, and to assure high quality of data. Our work addresses these gaps via targeted TLD crawling, specialised language discrimination between the mutually intelligible Croatian, Serbian, Bosnian and Montenegrin languages, and manual domain validation.
@@ 3.  CLASSLA-web Construction Pipeline
The goal of the crawling pipeline is to produce large, clean, and richly annotated web corpora. As shown in Figure 1, the process consists of 1) crawling based on top-level national domains and post-processing, 2) automatic text annotation with genre and topic information, and 3) linguistic annotation. In this section, we briefly present each step. For more details, refer also to the description of the CLASSLA-web crawling pipeline in Ljubešić and Kuzman (2024) and to the website that presents the CLASSLA-web collections (https://clarinsi.github.io/classla-web/).
@@
Following the methodology set up in the MaCoCu project (Bañón et al., 2022), we crawl national top-level domains (e.g., .si for Slovenian), as well as connected generic domains (.com and others). Crawling is performed with the MaCoCu crawler [^4] https://github.com/macocu/MaCoCu-crawler, which is based on the SpiderLing crawler (Suchomel et al., 2012) and designed for an efficient large-scale web text collection.
@@
Two language identification tools are used at the document and paragraph levels, namely a trigram classifier (Lui and Baldwin, 2014) and the Compact Language Detector 2 (CLD2, Sites, 2013) [^5] https://github.com/CLD2Owners/cld2. Language identification between Bosnian, Croatian, Montenegrin, and Serbian, comprising the HBS macro-language, has been shown to be particularly challenging, as these South Slavic languages exhibit a significant level of mutual intelligibility (Rupnik et al., 2023). We directly assign texts that come from national top-level domains to the corresponding web corpus, e.g., web domains from .hr are included in the Croatian web corpus. For generic top-level domains, we use an additional language identification tool specialised for distinguishing between Bosnian, Croatian, Montenegrin and Serbian, namely a Naive Bayes classifier trained on language-specific wordlists extracted from national TLDs (Rupnik et al., 2023). During the development of CLASSLA-web 2.0 corpora, we also explored language identification through zero-shot prompting with large language models, but the specialised classifier proved to be more reliable for HBS disambiguation.
@@
After crawling, we remove the boilerplate with the jusText [^6] http://corpus.tools/wiki/Justext tool (Pomikálek, 2011) and near-duplicates with the onion [^7] http://corpus.tools/wiki/Onion tool (Pomikálek, 2011), and discard exact duplicates. We discard texts that are not written in the target language, repair encoding, clean unwanted HTML elements, and filter overly short documents (comprising only short paragraphs with less than 70 characters or consisting of less than 75 words).
@@
In CLASSLA-web 2.0 corpora, a new annotation layer is introduced – in addition to genres, texts are annotated with topics to provide an additional insight into the corpora content. We use a multilingual news topic classifier (Kuzman and Ljubešić, 2025) covering 17 top level labels from the IPTC Media Topic NewsCodes schema (e.g., Politics, Science and Technology, Sport – for a detailed description of labels, refer to Kuzman and Ljubešić (2025)). On a manually-annotated test set in Croatian, Slovenian, Catalan, and Greek, the model achieves macro-F1 of 0.75 and micro-F1 of 0.73. For instances assigned with class probabilities below 0.6, we use the Mix label. As with genre annotation, the confidence threshold was identified based on manual analysis. We should note that the classifier has been evaluated for now only on news texts, as it was primarily intended for news topic classification. Thus, in Section 4.1, we analyse topics only within the News genre.
@@
All texts are linguistically annotated with the CLASSLA-Stanza pipeline (Ljubešić et al., 2024c) [^9] https://pypi.org/project/classla/, providing tokenization, lemmatization and morphosyntactic annotation with state-of-the-art accuracy for Slovenian [^10] Based on the SloBENCH benchmark at https://slobench.cjvt.si/leaderboard/view/11., and a high performance on Croatian, Serbian, Bulgarian, and Macedonian (Ljubešić et al., 2024c). We use the “web” module of the pipeline, which has been shown to perform well on heterogeneous web texts. The Croatian module is applied to Bosnian and Montenegrin. More detailed argumentation behind the module choices is provided in Ljubešić et al. (2024c).
@@
Although the overall CLASSLA-web pipeline is more or less the same between versions 1.0 and 2.0 of the CLASSLA-web corpora, three improvements were implemented during the development of the CLASSLA-web 2.0 version. First, near-duplicate removal was improved by masking numbers, punctuation, and links during paragraph-level deduplication to improve the efficiency of identifying texts that differ only in numbers. Second, we expanded the manual inspection of the top domains from only Slovenian and Croatian in CLASSLA-web 1.0 to all languages in CLASSLA-web 2.0. Third, we added topic annotation to complement genre labels.
@@
The CLASSLA-web 2.0 collection comprises all seven South Slavic national webs and languages – Bosnian (bs), Bulgarian (bg), Croatian (hr), Macedonian (mk), Montenegrin (cnr), Serbian (sr), and Slovenian (sl). In total, it consists of 17 billion words and 38 million texts, as reported in Table 1. The biggest corpus, namely the Bulgarian (CLASSLA-web.bg 2.0) corpus, comprises 6 billion words and 15 million texts. Other large corpora are Serbian (CLASSLA-web.sr 2.0), Croatian (CLASSLA-web.hr 2.0) and Slovenian (CLASSLA-web.sl 2.0) with 2.3 to 3.7 billion words and 5 to 7 million texts. Bosnian (CLASSLA-web.bs 2.0) and Macedonian (CLASSLA-web.mk 2.0) corpora are smaller, comprising 0.7 to 1 billion words and 2.1 to 2.5 million texts. The Montenegrin (CLASSLA-web.cnr 2.0) corpus is the smallest with 0.3 billion words and 0.8 million texts. However, as shown in Section 5.1, the new Montenegrin corpus is twice the size of the Montenegrin corpus included in the CLASSLA-web 1.0 collection and, to the best of our knowledge, represents the largest general Montenegrin corpus available.
@@ 4.1.  Genre and Topic Distribution
Genre and topic labels provide valuable insights into similarities as well as differences between web corpora. As shown in Figure 2, News content dominates in all national webs. This genre is especially dominant in Bosnian, Macedonian and Montenegrin corpora, where it exceeds 70%, whereas it represents 30–40% of texts in Croatian and Slovenian corpora. As in CLASSLA-web 1.0 corpora (Ljubešić and Kuzman, 2024), Promotion is more prevalent in Slovenian, Croatian and Bulgarian web corpora (12–20%) than in others. Another genre where there is a visible difference between corpora is Forum. Texts from forums seem to be much more frequent in Bulgarian, Serbian, and Slovenian webs (10–13%) than in Bosnian, Macedonian, and Montenegrin webs (below 5%).
@@
As News is the predominant genre in all web corpora and given that the topic classifier has been primarily validated on news texts, we analyse topics within the News genre. In Figure 3, we show the distribution of topics that are discussed in more than 10% of news texts in at least one corpus. Despite the fact that the news topic schema comprises 17 topic labels, only five categories appear with high enough frequency in the seven corpora, namely “Sport”, “Politics”, “Economy, Business and Finance”, “Arts, Culture, Entertainment and Media”, and “Human Interest”, in this order of average frequency across corpora. These five categories account for approximately 60% of news texts in each corpus. Sport is the most common topic of news texts in the majority of web corpora, and it even represents a fifth of all news in the Macedonian web corpus. In Bulgarian and Montenegrin web corpora, the most frequent topic of news texts is Politics, accounting for 20% of news texts within the Montenegrin corpus. The distribution of the other three most frequent topic categories is relatively consistent across the South Slavic corpora, except in the case of Arts, Culture, Entertainment and Media which is notably more prevalent in the Croatian corpus compared to the others, representing almost 15% of Croatian news texts.
@@
To understand how national webs evolve over short time spans, we compare the CLASSLA-web 2.0 corpora with CLASSLA-web 1.0 (Ljubešić and Kuzman, 2024). Both web corpus collections were compiled using an identical construction pipeline (see Section 3), with the 1.0 version having been collected from the web two to three years prior to the 2.0 version, specifically in 2021 and 2022. In the following sections, we examine the size difference between the two corpora collections (Section 5.1), content overlap (Section 5.2), and the presence of automatically generated texts and other low quality content (Section 5.3).
@@ 5.1.  Size Gains
As shown in Figure 4, the second crawling iteration produced larger web corpora, showing the increase of content on the web. Compared to the size of the CLASSLA-web 1.0 corpora, the web corpora in the 2.0 corpus collection comprise 57% more words and 46% more texts, that is, roughly half as many more words and texts than in the first version. Interestingly, web growth is uneven across South Slavic webs – the Slovenian and Croatian webs exhibit modest increase of size, with a 30% change in size in terms of number of words, while the size of Bulgarian and Montenegrin web corpora nearly doubled in the second crawling iteration. These differences in sizes reflect the high growth of content on the web rather than any differences in the crawling methodology, as both the CLASSLA-web 1.0 and CLASSLA-web 2.0 corpora have been produced with practically the same pipeline.